Determine the direction and magnitude of the electric field at point P. The charges are 

separated by a distance 2𝑎, and point P is a distance 𝑥 from the midpoint between the two 

charges. Express your answer in terms of 𝑄, 𝑥, 𝑎, and 𝑘.



1
Expert's answer
2021-05-27T18:39:57-0400

Electric field, "E=\frac{kQ}{x^2}"

let charge to the right=+Q and charge to the left =-Q

"E_1=\frac{-kQ}{(x-a)^2}"

"E_2=\frac{kQ}{(x+a)^2}"

net electric field, "E=E_1+E_2=" "\frac{-kQ}{(x-a)^2}+""\frac{kQ}{(x+a)^2}"

"E=kQ\frac{[(x^2+a^2-2ax)-(x^2+a^2+2ax)]}{(x^2-a^2)^2}"

"E=\frac{-4kQax}{(x^2-a^2)^2}"